two angles of a triange have the same measure and the third angle is 30 degrees greater than the other two. Find the measures of each angle.
Found 2 solutions by edjones, RAY100:
Answer by edjones(8007)

Let the 2 equal angles be x.
2x+x+30=180
3x=150
x=50
x+30=80
.
Ed

Answer by RAY100(1637)

remember the sum if angles of a triangle =180 deg
gives angles of,, x,x,(x+30)
sum = 3x +30=180
subt 30 both sides
3x=180-30=150
divide both sides by 3
x=50
answer, sides of 50,50,80
check 50+50+80=180 ok
80-50=30 ok